OpenStudy (anonymous):

When a strip of zinc is placed in a solution of silver nitrate, Zn replaces Ag to form zinc nitrate, and metallic silver is formed. How many moles of silver will be generated if 1.30 moles of zinc is placed into the silver nitrate solution? 2AgNO3+Zn-> 2Ag+Zn(NO3)2 A) 0.65 moles B) 1.30 moles C) 2.60 moles D) 6.20 moles

OpenStudy (aaronq):

you need the ratio of their moles and coefficients, like this: \(\sf \dfrac{moles~of~Zn}{Zn's~coefficient}=\dfrac{moles~of~Ag}{Ag's~coefficient}\) plug in what you know and solve it (like algebra) For example: \(\dfrac{2}{3}=\dfrac{x}{2}\rightarrow x=\dfrac{2*2}{3}\)